# maximum a posteriori

### v1

@@ B.1 Bias and Variance of BOLT
Let $c(x)\vcentcolon=q_{1}p_{1}(x)-q_{2}p_{2}(x)$. A pointwise maximizer of c(x)h(x)c(x)\,h(x) over $h(x)\in[-1,0]$ is
$h^{\star}(x)\;=\;\begin{cases}-1,&U(x)<\tau,\\[2.0pt] 0,&U(x)\geq\tau,\end{cases}$ (79)
and $\mathcal{L}_{\mathrm{BOLT}}(h^{\star})=\varepsilon_{\mathrm{Bayes}}$.
@@
We start by showing the endpoint optimality. For fixed $x$, the map $h\mapsto c(x)h$ is affine on [1,0][-1,0], hence maximized at an endpoint chosen by the sign of $c(x)$, i.e., $h^{\star}(x)=-1$ if $c(x)<0$ and $h^{\star}(x)=0$ otherwise. Since $c(x)<0$ holds if and only if $U(x)<\tau$, we can conclude (79).
We next show the tightness of the result. Using the definition of $\mathcal{L}_{\mathrm{BOLT}}$ (Def. 5, Sec. S1) and substituting $h^{\star}$, we have
$\mathcal{L}_{\mathrm{BOLT}}(h^{\star})$ $=q_{2}\;-\;\Big(q_{1}\,\mathbb{E}_{X\sim P_{1}}[h^{\star}(X)]-q_{2}\,\mathbb{E}_{X\sim P_{2}}[h^{\star}(X)]\Big)$ (80)
$=q_{2}\;-\;\int\big(q_{1}p_{1}(x)-q_{2}p_{2}(x)\big)\,h^{\star}(x)\,dx$ (81)
$=q_{2}\;-\;\int\big(q_{2}p_{2}(x)-q_{1}p_{1}(x)\big)\,1\{U(x)<\tau\}\,dx$ (82)
$=q_{2}\;-\;\int\big[\,q_{2}p_{2}(x)-q_{1}p_{1}(x)\,\big]_{+}\,dx$ (83)
@@
Here, we have used the fact that a $1\{a>0\}=[a]_{+} a$nd $q_{2}=\int q_{2}p_{2}$. This completes the proof. ∎
@@
Corollary 3 (Bayes error as a hinge of likelihood ratio).
With $h^{\star}$ in (79), define the hinge map as
@@
The Bayes error rate is then given by
@@
Starting from (81) and using $h^{\star}=-1\{U<\tau\}$,
$q_{1}\mathbb{E}_{P_{1}}[h^{\star}]-q_{2}\mathbb{E}_{P_{2}}[h^{\star}]$ $=\int\!\big(q_{1}p_{1}-q_{2}p_{2}\big)h^{\star}\,dx=\int\!\big(q_{2}p_{2}-q_{1}p_{1}\big)\,1\{U<\tau\}\,dxlabel{eq:S3-swap}$ (87)
$=\int\!\big[\,q_{2}p_{2}-q_{1}p_{1}\,\big]_{+}\,dx=\mathbb{E}_{X\sim P_{2}}\!\big[\,q_{2}-q_{1}U(X)\,\big]_{+},$ (88)
where the last equality is obtain from dividing by as $p_{2}(x)$ on $\{p_{2}>0\}$ (the complement contributes 0). Plugging into the expression for $\mathcal{L}_{\mathrm{BOLT}}(h^{\star})$, we obtain (86). ∎
@@
Suppose we observe $M_{1}$ samples from class $C_{1}$ and $M_{2}$ from class $C_{2}$ (with total $M=M_{1}+M_{2}$), and set empirical priors $\widehat{q}_{i}\vcentcolon=M_{i}/M$. Let $\widehat{U}$ estimate $U$, and let $X^{(2)}_{1},\ldots,X^{(2)}_{M_{2}}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}P_{2}$ be the $M_{2}$ samples of class-$C_{2}$ used to estimate the expectation over $P_{2}$ in (86). Define
@@
Note $0\leq t_{0}(u)\leq q_{2}\leq 1$ and $0\leq\widehat{t}_{0}(u)\leq\widehat{q}_{2}\leq 1$, so all summands are bounded in $[0,1]$.
@@
Assume i.i.d. draws and bounded outputs in $[0,1]$. Then
$\big|\,\mathbb{E}[\widehat{\varepsilon}_{\mathrm{BOLT}}]-\varepsilon_{\mathrm{Bayes}}\,\big|$ $\;\leq\;q_{1}\,\mathbb{E}\!\big[\,|\widehat{U}(X)-U(X)|\,\big]\;+\;O(M^{-1/2}),$ (91)
@@
In particular, if $\mathbb{E}[\,|\widehat{U}(X)-U(X)|\,]\leq\varepsilon_{0}$, then
@@
Bias. By (86)-(89) and adding and subtracting $\mathbb{E}_{X\sim P_{2}}\!\big[t_{0}(\widehat{U}(X))\big]$, we have:
$\mathbb{E}[\widehat{\varepsilon}_{\mathrm{BOLT}}]-\varepsilon_{\mathrm{Bayes}}=\underbrace{\mathbb{E}_{X\sim P_{2}}\!\big[t_{0}(U(X))-t_{0}(\widehat{U}(X))\big]}_{\text{plug-in error}}+\underbrace{\mathbb{E}\!\Big[t_{0}(\widehat{U}(X))-\tfrac{1}{M_{2}}\sum_{i=1}^{M_{2}}\widehat{t}_{0}\big(\widehat{U}(X_{i})\big)\Big]}_{\text{sampling/priors}}.$ (94)
By Lemma 5, $t_{0}$ is $q_{1}$-Lipschitz, so $\big|\mathbb{E}_{P_{2}}[t_{0}(U)-t_{0}(\widehat{U})]\big|\leq q_{1}\,\mathbb{E}_{P_{2}}[|\widehat{U}-U|]$. For the sampling term, write $Z_{i}=\widehat{t}_{0}(\widehat{U}(X_{i}))\in[0,1]$; Hoeffding gives $\mathbb{E}\big|\tfrac{1}{M_{2}}\sum Z_{i}-\mathbb{E}Z_{1}\big|=O(M_{2}^{-1/2})$. The empirical priors $\widehat{q}_{i}=M_{i}/M$ add a zero-mean fluctuation with variance $O(M^{-1})$, hence $O(M^{-1/2})$ in expectation. With fixed class proportions ($M_{2}=\Theta(M)$), these contributions are of the same order, yielding (91).
Variance.  Using $\widehat{\varepsilon}_{\mathrm{BOLT}}=\widehat{q}_{2}-\tfrac{1}{M_{2}}\sum_{i=1}^{M_{2}}Z_{i}$ with bounded $Z_{i}$, $\mathrm{Var}\!\big(\tfrac{1}{M_{2}}\sum_{i}Z_{i}\big)=O(M_{2}^{-1})$ and $\mathrm{Var}(\widehat{q}_{2})=O(M^{-1})$. Any covariance between these two bounded averages is $O(M^{-1})$ and is dominated by the same order; hence (92). ∎
@@
where $W$ is a width/capacity proxy, $N$ is the number of training samples used to learn $\widehat{U}$, and $\gamma>0$ depends on the target smoothness and the approximating family (Lu et al.,, 2020; Shalev-Shwartz and Ben-David,, 2014).
@@ Appendix C Further Results on BOLT-GAN
$D^{(\pi)}_{\mathrm{Lip}}(g)$ $\;\vcentcolon=\;\sup_{h\in\mathcal{H}_{\mathrm{Lip}}}\mathbb{E}_{X\sim P_{\mathrm{data}}}\!\big[h(X)\big]-(1-\pi)\,\mathbb{E}_{X\sim P_{g}}\!\big[h(X)\big],$ (98)
@@
The case $\pi>\tfrac{1}{2}$ follows by exchanging labels (replacing $\pi$ with $1-\pi$, i.e., swaping $P_{\mathrm{data}}$ and $P_{g}$). Thus the conclusions above extend to all $\pi\in(0,1]$ by symmetry.
@@ Appendix D Implementation of Lipschitz BOLT-GAN
This section records the implementation details for our Lipschitz BOLT–GAN experiments. We explain the properties we rely on and the techniques used to enforce the $1$-Lipschitz constraint on the discriminator (critic). As discussed in the main paper, Lipschitz continuity is key both (i) to recover meaningful dual objectives (e.g., the KR dual for $W_{1}$) and (ii) to stabilize adversarial training; see the TV/Wasserstein connections in Secs. 3.2–3.3. In our code, the gradient penalty is applied to the raw critic output $\tilde{h}_{\theta}$ (pre‑activation), and the bounded score $h_{\theta}=\sigma(\tilde{h}_{\theta})\in[0,1]$ is used inside the BOLT objective, as summarized in Algorithm 1. (Gulrajani et al.,, 2017; Miyato et al.,, 2018, Sec. 3).
@@ D.1 Preliminaries on Lipschitz continuity
Lemma 9 (Basic Lipschitz calculus (Miyato et al.,, 2018)).
@@
Lemma 10 (Layerwise Lipschitz constants in common architectures (Miyato et al.,, 2018; Anil et al.,, 2019)).
@@ D.2 Spectral Normalization (SN)
Method. Each linear layer $W$ is rescaled by an estimate of its top singular value $\sigma(W)$ via power iteration with persistent vectors $(u,v)$: $\bar{W}\leftarrow W/\sigma(W)$. For convolutions, $\sigma(W)$ is the operator norm of the induced linear map and can be estimated by alternating conv/transpose‑conv passes. (Miyato et al.,, 2018).
Guarantee. If every linear/convolutional layer is normalized so that $\|W\|_{2}\leq 1$ and all activations are $1$-Lipschitz, then by Lemmas 910 the whole network is $1$-Lipschitz (up to residual additions; see “Residual blocks” below). This yields strong stability at low computational cost and is widely adopted in GAN critics. (Miyato et al.,, 2018).
@@ D.3 Gradient Penalty (GP)
Method. Gradient penalty encourages $\|\nabla_{x}\tilde{h}_{\theta}(x)\|_{2}\approx 1$ on a chosen sampling distribution. The most common choice is the interpolation penalty (Gulrajani et al.,, 2017): draw $(x_{\mathrm{real}},x_{\mathrm{fake}})$, form $\hat{x}=\alpha x_{\mathrm{real}}+(1-\alpha)x_{\mathrm{fake}}$ with $\alpha\sim\mathrm{Unif}[0,1]$, and add
@@
Guarantee (local, not global). The penalty (104) is a soft constraint: it is zero iff the gradient norm equals $1$ on the support where it is evaluated (here, segments between real/fake samples). It does not by itself prove a global $1$-Lipschitz bound, but it aligns with the KR optimal‑critic condition and markedly improves training stability. (Gulrajani et al.,, 2017).
@@ D.4 Weight clipping
Method. Clip each parameter after every update: $w\leftarrow\mathrm{clip}(w,\,-\mathcal{\mathcal{}}{L}_{\mathrm{wc}},\,\mathcal{\mathcal{}}{L}_{\mathrm{wc}})$ (Arjovsky et al.,, 2017). This loosely bounds per‑layer operator norms via element‑wise bounds.
@@
so a depth‑$L$ network with 11-Lipschitz activations satisfies $\mathrm{Lip}(f)\leq\prod_{\ell=1}^{L}\sqrt{m_{\ell}n_{\ell}}\,\mathcal{\mathcal{}}{L}_{\mathrm{wc}}$. This yields at best a coarse global bound and often harms capacity. In practice we prefer SN or GP. (Arjovsky et al.,, 2017; Miyato et al.,, 2018).

### v2

@@ B.1 Bias and Variance of BOLT
Assume $q_{1},q_{2}>0$ and that $P_{1},P_{2}$ admit densities $p_{1},p_{2}$ w.r.t. a common dominating measure. Define
$U(x)\coloneqq\frac{p_{1}(x)}{p_{2}(x)}\quad\text{on }\{p_{2}(x)>0\},\qquad\tau\coloneqq\frac{q_{2}}{q_{1}},$ (86)
and let
$c(x)\vcentcolon=q_{1}p_{1}(x)-q_{2}p_{2}(x).$ (87)
A pointwise maximizer of c(x)h(x)c(x)\,h(x) over $h(x)\in[0,1]$ is
$h^{\star}(x)\;=\;\begin{cases}0,&U(x)<\tau,\\[2.0pt] 1,&U(x)\geq\tau,\end{cases}$ (88)
and the BOLT bound is tight:
$\mathcal{L}_{\mathrm{BOLT}}(h^{\star})=\varepsilon_{\mathrm{Bayes}},$ (89)
where (consistent with Def. 6 )
$\mathcal{L}_{\mathrm{BOLT}}(h)\;\coloneqq\;q_{1}\;+\;q_{2}\,\mathbb{E}_{X\sim P_{2}}[h(X)]\;-\;q_{1}\,\mathbb{E}_{X\sim P_{1}}[h(X)].$ (90)
@@
Endpoint optimality. For fixed $x$, the map $h\mapsto c(x)h$ is affine on $[0,1]$, hence maximized at an endpoint chosen by the sign of $c(x)$: $h^{\star}(x)=1$ if $c(x)\geq 0$ and $h^{\star}(x)=0$ if $c(x)<0$. Since $c(x)<0$ holds if and only if $q_{1}p_{1}(x)<q_{2}p_{2}(x)$, i.e. $U(x)<\tau$, we obtain (88).
Tightness. Using the definition (90) and substituting $h^{\star}$, we have
$\mathcal{L}_{\mathrm{BOLT}}(h^{\star})$ $=q_{1}\;+\;q_{2}\,\mathbb{E}_{X\sim P_{2}}[h^{\star}(X)]\;-\;q_{1}\,\mathbb{E}_{X\sim P_{1}}[h^{\star}(X)]$ (91)
$=q_{1}\;+\;\int\big(q_{2}p_{2}(x)-q_{1}p_{1}(x)\big)\,h^{\star}(x)\,dx$ (92)
$=q_{1}\;+\;\int\big(q_{2}p_{2}(x)-q_{1}p_{1}(x)\big)\,\mathbf{1}\{U(x)\geq\tau\}\,dx$ (93)
$=q_{1}\;-\;\int\big(q_{1}p_{1}(x)-q_{2}p_{2}(x)\big)\,\mathbf{1}\{U(x)\geq\tau\}\,dx$ (94)
$=q_{1}\;-\;\int\big[\,q_{1}p_{1}(x)-q_{2}p_{2}(x)\,\big]_{+}\,dx$ (95)
$=q_{1}\;-\;\int\Big(\max\{q_{1}p_{1}(x),q_{2}p_{2}(x)\}-q_{2}p_{2}(x)\Big)\,dx$ (96)
$=q_{1}\;-\;\int\max\{q_{1}p_{1}(x),q_{2}p_{2}(x)\}\,dx\;+\;\int q_{2}p_{2}(x)\,dx$ (97)
@@
Here we used [a]+=a $_{+}=a\,\mathbf{1}\{a>0\}$ and the identity $b-a]_{+}=\max\{b,a\}-a$ (with $b=q_{1}p_{1}(x)$, $a=q_{2}p_{2}(x)$), and also $\int q_{2}p_{2}(x)\,dx=q_{2}$. This completes the proof. ∎
@@
Corollary 4 (Bayes error as a hinge of the likelihood ratio).
Define the hinge map
@@
Then the Bayes error rate is given by
@@
Recall
$\varepsilon_{\mathrm{Bayes}}=1-\int\max\{q_{1}p_{1}(x),q_{2}p_{2}(x)\}\,dx=\int\min\{q_{1}p_{1}(x),q_{2}p_{2}(x)\}\,dx.$ (101)
Using $\min\{a,b\}=b-[\,b-a\,]_{+}$ with $a=q_{1}p_{1}(x)$ and $b=q_{2}p_{2}(x)$, we get
$\varepsilon_{\mathrm{Bayes}}$ $=\int q_{2}p_{2}(x)\,dx-\int[\,q_{2}p_{2}(x)-q_{1}p_{1}(x)\,]_{+}\,dx$
$=q_{2}-\int p_{2}(x)\,[\,q_{2}-q_{1}U(x)\,]_{+}\,dx=q_{2}-\mathbb{E}_{X\sim P_{2}}\!\Big[[\,q_{2}-q_{1}U(X)\,]_{+}\Big],$ (102)
which is (100). ∎
@@
Suppose we observe $M_{1}$ samples from class $C_{1}$ and $M_{2}$ from class $C_{2}$ (with total $M=M_{1}+M_{2}$), and set empirical priors
$\widehat{q}_{i}\vcentcolon=\frac{M_{i}}{M},\qquad i\in\{1,2\}.$ (103)
Let $\widehat{U}$ estimate $U$, and let $X^{(2)}_{1},\ldots,X^{(2)}_{M_{2}}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}P_{2}$ be the $M_{2}$ samples of class-$C_{2}$ used to estimate the expectation over $P_{2}$ in (100). Define
@@
Note that $0\leq t_{0}(u)\leq q_{2}\leq 1$ and $0\leq\widehat{t}_{0}(u)\leq\widehat{q}_{2}\leq 1$, so all summands are bounded in $[0,1]$.
@@
Assume i.i.d. draws and bounded hinge summands in $[0,1]$. Then
$\big|\,\mathbb{E}[\widehat{\varepsilon}_{\mathrm{BOLT}}]-\varepsilon_{\mathrm{Bayes}}\,\big|$ $\;\leq\;q_{1}\,\mathbb{E}_{X\sim P_{2}}\!\big[\,|\widehat{U}(X)-U(X)|\,\big]\;+\;O(M^{-1/2}),$ (106)
@@
In particular, if $\mathbb{E}_{X\sim P_{2}}\!\big[\,|\widehat{U}(X)-U(X)|\,\big]\leq\varepsilon_{0}$, then
@@
Bias. By (104) and adding/subtracting $\mathbb{E}_{X\sim P_{2}}\!\big[t_{0}(\widehat{U}(X))\big]$, we have
$\mathbb{E}[\widehat{\varepsilon}_{\mathrm{BOLT}}]-\varepsilon_{\mathrm{Bayes}}$ $=\underbrace{\mathbb{E}_{X\sim P_{2}}\!\big[t_{0}(U(X))-t_{0}(\widehat{U}(X))\big]}_{\text{plug-in error}}+\underbrace{\mathbb{E}\!\Bigg[\mathbb{E}_{X\sim P_{2}}\!\big[t_{0}(\widehat{U}(X))\big]-\frac{1}{M_{2}}\sum_{i=1}^{M_{2}}\widehat{t}_{0}\big(\widehat{U}(X^{(2)}_{i})\big)\Bigg]}_{\text{sampling/priors}}.$ (109)
By Lemma 6, $t_{0}$ is $q_{1}$-Lipschitz, so
$\Big|\mathbb{E}_{X\sim P_{2}}\!\big[t_{0}(U(X))-t_{0}(\widehat{U}(X))\big]\Big|\leq q_{1}\,\mathbb{E}_{X\sim P_{2}}\!\big[|\widehat{U}(X)-U(X)|\big].$ (110)
For the sampling/priors term, write $Z_{i}=\widehat{t}_{0}(\widehat{U}(X^{(2)}_{i}))\in[0,1]$. Hoeffding implies $\mathbb{E}\big|\tfrac{1}{M_{2}}\sum_{i=1}^{M_{2}}Z_{i}-\mathbb{E}Z_{1}\big|=O(M_{2}^{-1/2})$. The empirical priors $\widehat{q}_{i}=M_{i}/M$ add a zero-mean fluctuation with variance $O(M^{-1})$, hence $O(M^{-1/2})$ in expectation. With fixed class proportions ($M_{2}=\Theta(M)$), these contributions yield (106).
Variance. Using $\widehat{\varepsilon}_{\mathrm{BOLT}}=\widehat{q}_{2}-\tfrac{1}{M_{2}}\sum_{i=1}^{M_{2}}Z_{i}$ with bounded $Z_{i}$, $\mathrm{Var}\!\big(\tfrac{1}{M_{2}}\sum_{i}Z_{i}\big)=O(M_{2}^{-1})$ and $\mathrm{Var}(\widehat{q}_{2})=O(M^{-1})$. Any covariance between these two bounded averages is $O(M^{-1})$, hence (107). ∎
@@
where $W$ is a width/capacity proxy, $N$ is the number of training samples used to learn $\widehat{U}$, and $\gamma>0$ depends on the target smoothness and the approximating family (Lu et al., 2020; Shalev-Shwartz and Ben-David, 2014).
@@ Appendix C Further Results on BOLT-GAN
$D^{(\pi)}_{\mathrm{Lip}}(g)$ $\;\vcentcolon=\;\pi\sup_{h\in\mathcal{H}_{\mathrm{Lip}}}\mathbb{E}_{X\sim P_{\mathrm{data}}}\!\big[h(X)\big]-(1-\pi)\,\mathbb{E}_{X\sim P_{g}}\!\big[h(X)\big],$ (114)
@@
The case $\pi>\tfrac{1}{2}$ follows by exchanging labels (replacing $\pi$ with $1-\pi$, i.e., swapping $P_{\mathrm{data}}$ and $P_{g}$). Thus the conclusions above extend to all $\pi\in(0,1]$ by symmetry.
@@ Appendix D Implementation of Lipschitz BOLT-GAN
This section records the implementation details for our Lipschitz BOLT–GAN experiments. We explain the properties we rely on and the techniques used to enforce the $1$-Lipschitz constraint on the discriminator (critic). As discussed in the main paper, Lipschitz continuity is key both (i) to recover meaningful dual objectives (e.g., the KR dual for $W_{1}$) and (ii) to stabilize adversarial training; see the TV/Wasserstein connections in Secs. 3.2–3.3. In our code, the gradient penalty is applied to the raw critic output $\tilde{h}_{\theta}$ (pre‑activation), and the bounded score $h_{\theta}=\sigma(\tilde{h}_{\theta})\in[0,1]$ is used inside the BOLT objective, as summarized in Algorithm 1. (Gulrajani et al., 2017; Miyato et al., 2018, Sec. 3).
@@ D.1 Preliminaries on Lipschitz continuity
Lemma 10 (Basic Lipschitz calculus (Miyato et al., 2018)).
@@
Lemma 11 (Layerwise Lipschitz constants in common architectures (Miyato et al., 2018; Anil et al., 2019)).
@@ D.2 Spectral Normalization (SN)
Method. Each linear layer $W$ is rescaled by an estimate of its top singular value $\sigma(W)$ via power iteration with persistent vectors $(u,v)$: $\bar{W}\leftarrow W/\sigma(W)$. For convolutions, $\sigma(W)$ is the operator norm of the induced linear map and can be estimated by alternating conv/transpose‑conv passes. (Miyato et al., 2018).
Guarantee. If every linear/convolutional layer is normalized so that $\|W\|_{2}\leq 1$ and all activations are $1$-Lipschitz, then by Lemmas 1011 the whole network is $1$-Lipschitz (up to residual additions; see “Residual blocks” below). This yields strong stability at low computational cost and is widely adopted in GAN critics. (Miyato et al., 2018).
@@ D.3 Gradient Penalty (GP)
Method. Gradient penalty encourages $\|\nabla_{x}\tilde{h}_{\theta}(x)\|_{2}\approx 1$ on a chosen sampling distribution. The most common choice is the interpolation penalty (Gulrajani et al., 2017): draw $(x_{\mathrm{real}},x_{\mathrm{fake}})$, form $\hat{x}=\alpha x_{\mathrm{real}}+(1-\alpha)x_{\mathrm{fake}}$ with $\alpha\sim\mathrm{Unif}[0,1]$, and add
@@
Guarantee (local, not global). The penalty (120) is a soft constraint: it is zero iff the gradient norm equals $1$ on the support where it is evaluated (here, segments between real/fake samples). It does not by itself prove a global $1$-Lipschitz bound, but it aligns with the KR optimal‑critic condition and markedly improves training stability. (Gulrajani et al., 2017).
@@ D.4 Weight clipping
Method. Clip each parameter after every update: $w\leftarrow\mathrm{clip}(w,\,-\mathcal{\mathcal{}}{L}_{\mathrm{wc}},\,\mathcal{\mathcal{}}{L}_{\mathrm{wc}})$ (Arjovsky et al., 2017). This loosely bounds per‑layer operator norms via element‑wise bounds.
@@
so a depth‑$L$ network with 11-Lipschitz activations satisfies $\mathrm{Lip}(f)\leq\prod_{\ell=1}^{L}\sqrt{m_{\ell}n_{\ell}}\,\mathcal{\mathcal{}}{L}_{\mathrm{wc}}$. This yields at best a coarse global bound and often harms capacity. In practice we prefer SN or GP. (Arjovsky et al., 2017; Miyato et al., 2018).